### v1

@@ 5.2.1 Impact of Stages in PPT Method
Figure 4: An evaluation of the effectiveness of the PPT strategy on mathematical reasoning tasks. We present the zero-shot Pass@1 results on the MATH and GSM8k benchmarks, across three cumulative stages of the PPT strategy based on the DeepSeek-Math and Llama-3.1 models. . The results highlight the cumulative effectiveness of the PPT strategy, showing increased performance for both models across all benchmark datasets with each progressive stage. These results indicate that the PPT strategy is a viable method to enhance mathematical reasoning capabilities.
To evaluate the effectiveness of PPT methods, we employ two base models, Llama-3.1-8B (Dubey et al., 2024) and Deepseek-Math-7B-Base (Shao et al., 2024). We present the results of the PPT method across different stages on these two models for the MATH and GSM8K benchmarks. As shown in Figure 4, the vanilla models exhibit limited performance on both benchmarks. After stage ①, the performance of all models improves significantly. For instance, Llama-3.1-Base achieves absolute gains of $47.9\%$ on MATH and $61.0\%$ on GSM8K. This demonstrates the necessity of understanding mathematical concepts and warming up the models. Subsequently, we conduct stage ② training, which utilizes data with NLR and AR. The limited performance gains are potential because the pre-training data already includes code and natural language corpus. Both stage ① and stage ② primarily reactivate the base models in the mathematical domain, and stage ① already incorporates NLR training. In the final stage ③, we consider training with three paradigms, and all models experience performance enhancements. This suggests that introducing rare or unseen paradigms helps to improve the reasoning ability of learned paradigms.
5.2.2 Impact of Reasoning Paradigms Order
We investigated the impact of varying the sequence of reasoning paradigms employed by CoR-Math-7B during zero-shot inference for arithmetic computation tasks. Our experiments involved modifying only the prompt structure to enforce different paradigm orders. Given that NLR closely aligns with the pre-training of LLMs, it was consistently positioned as the initial paradigm in our experiments. Table 4 presents the accuracy for the MATH and GSM8K benchmarks under two different paradigm orders. The results indicate that performing SR before AR achieves the best performance. Specifically, on the MATH benchmark, this order achieves an accuracy of $66.7\%$, an improvement over the $49.9\%$ attained by the sequence that AR before SR. A possible explanation for this observation lies in the synergistic effect of the ordered paradigms. SR plays a crucial role in decomposing the problem into manageable sub-steps, thereby providing a structured foundation for subsequent AR. By placing SR before AR, we hypothesize that the model is better equipped to perform accurate calculations, leveraging the intermediate reasoning steps generated by SR. Furthermore, positioning AR immediately preceding the Summary phase may facilitate a more coherent integration of the computational outcome into the final answer.
@@ 5.2.2 Impact of Reasoning Paradigms Order
[…]
MATH 66.7 49.9
GSM8K 88.7 84.2
5.2.3 Impact of Scaling Model Size
@@
Table 5: Zero-shot Pass@1 results on the MATH and GSM8k benchmarks across different model scales. The best scores are in bold.
Model MATH GSM8k
Qwen2.5-Math-1.5B (CoR) 57.6 84.5
Qwen2.5-Math-7B (CoR) 64.7 90.0
Table 5 investigates the performance of various base models with different parameters under the CoR framework. We employ the Qwen2.5-Math models (Yang et al., 2024) with $1.5$ billion and $7$ billion parameters as the base model. Results demonstrate that our CoR framework achieves superior performance with increasing parameter scale. Specifically, when comparing a model with $7$ billion parameters to one with $1.5$ billion parameters, the former attained an absolute increase of $7.1\%$ on the MATH benchmark and $5.5\%$ on the GSM8K benchmark. This evidence supports the claim that CoR exhibits scalability with respect to model parameter scaling.
6 Discussion on Reasoning Hierarchy: Paradigms, Paths, and Steps
@@
Figure 5: An overview of the reasoning hierarchical structure. (a) A single reasoning paradigm depicts multiple distinct reasoning paths. (b) An example of single-paradigm reasoning includes one reasoning path, which contains several reasoning steps. (c) Multi-paradigm reasoning includes several distinct reasoning paradigms.
As shown in Figure 5, this paper posits that reasoning texts generated by LLMs exhibit a reasoning hierarchical structure, which consists of three levels: reasoning paradigms, reasoning paths, and reasoning steps.
@@
  • Reasoning paradigms comprise one or more reasoning paths. They often contain multiple potential final answers, thus necessitating a summarization method, such as a summary module, to extract the ultimate answer. Furthermore, a reasoning paradigm utilizes a single knowledge media, such as natural language.
Figure 6: A comparison of reasoning paradigms in advanced studies. (a) deep reasoning, characterized by serial concatenation of reasoning paths. (b) interleaved reasoning, which integrates secondary methods for guidance within a dominant paradigm. (c) multi-paradigm reasoning (Our proposed CoR), a framework designed to harness the synergistic effects between different reasoning paradigms.
@@
Furthermore, current studies can be categorized based on their focus. As shown in Figure 5 (a), contemporary work concentrates within a single paradigm, optimizing along two dimensions: depth (the number of reasoning steps) and width (the number of reasoning paths). For instance, regarding reasoning depth, the CoT method (Wei et al., 2022) employs prompts to increase intermediate steps within one reasoning path to achieve higher performance. Concerning reasoning width, some approaches (Wang et al., 2023) involve altering sampling techniques to generate multiple distinct reasoning paths. Random sampling exemplifies this. Other studies (Zhu et al., 2023) employ scoring mechanisms for reasoning steps, guiding LLMs to generate several more desirable reasoning paths. Monte Carlo search illustrates this. Building upon these generated reasoning paths, existing work proposes various integration strategies to obtain the final answer, such as Best-of-N and Self-consistency (Wang et al., 2023). Specifically, as shown in Figure 6 (a), recent advanced studies introduce the deep reasoning method, which focuses on generating serial concatenation of reasoning paths followed by summarization (like OpenAI o1). Since these methods are based on different optimization dimensions, we can combine them feasibly in applications.
@@
Specifically, some approaches prioritize a dominant paradigm for reasoning yet hope to integrate other paradigms for guidance. For instance, as shown in Figure 6 (b), the reasoning process of ToRA (Gou et al., 2024) presents an interleaved reasoning method with different knowledge media. It involves initial natural language generation, which is followed by code generation. Subsequently, the process awaits the results of code execution before generating further natural language. InternLM2.5-StepProver (Wu et al., 2024) exhibits a similar pattern. It incorporates natural language annotations to support symbolic reasoning steps. However, these interleaved reasoning approaches do not constitute true multi-paradigm reasoning. The primary paradigm can independently achieve the final answer without the supplementary paradigms. Consequently, these methods imply a single-paradigm reasoning approach enhanced by other paradigms, rather than genuine multi-paradigm reasoning.
@@
Current methods often overlook the synergistic effects between paradigms and underestimate the importance of complete reasoning for achieving accurate results. To address this challenge, as shown in Figure 5 (c) and Figure 6 (c), we introduce Chain-of-Reasoning (CoR), a unified reasoning framework capable of multi-paradigm reasoning. This framework embodies the following potential advantages:
@@
Sequential reasoning dependency. As a type of multi-paradigm reasoning, CoR is not a mere accumulation of isolated steps; instead, it represents a coherent, interconnected process. The output from an earlier paradigm functions not only as input for a subsequent paradigm but also informs the foundational information for its reasoning. For instance, knowledge derived from a natural language paradigm guides subsequent algorithmic paradigms, thereby enhancing the efficiency and accuracy of code generation based on detailed natural language reasoning. This mechanism also provides substantial context from prior paradigms to later paradigms and operates as a form of preliminary rehearsal where subsequent paradigms can follow correct reasoning steps or rectify incorrect ones.
@@
7 Conclusions
This paper introduces Chain-of-Reasoning (CoR), a novel unified reasoning framework that synergistically combines natural language, algorithmic, and symbolic reasoning to solve mathematical problems. Our approach demonstrates the potential of integrating diverse reasoning paradigms in the reasoning process, enabling language models to tackle complex mathematical challenges effectively. CoR achieves significant performance gains on challenging mathematical reasoning tasks, substantially outperforming existing single-paradigm approaches on benchmarks like MATH and MiniF2F. In addition, CoR introduces a new, orthogonal perspective to traditional test-time scaling by leveraging multi-paradigm reasoning paths to enhance computational efficiency and performance. Furthermore, our proposed CoR framework offers valuable insights for future research, contributing to the development of more robust and versatile reasoning agents capable of leveraging the strengths of multiple reasoning modalities.
Appendix
Appendix A Experiment Details
A.1 The detail of the Universal Text Template
@@
As shown in Figure 7, we apply the designed universal text template on all training samples.
A.2 Example Prompts for Dataset Enhancement
@@
Figure 9 and Figure 8 provide examples of such prompts designed for augmenting and refining formal proofs within the Lean 4 theorem prover for the Numina-TIR and Lean-Workbook datasets, respectively. The prompt specifies the input and output format and includes placeholders for the problem statement, informal proof, and corresponding formal proof in Lean 4. It also incorporates $N_{\text{shot}}$ few-shot examples to further guide the model. In our experiments, consistent with the common settings employed in (Jiang et al., 2023; Yang et al., 2024), we use $N_{\text{shot}}=8$. This structured approach and setting are adapted and applied to all datasets to ensure comprehensive coverage and logical consistency across the enhanced dataset.
@@
A.3 Details of Training Settings
In all stages of the PPT method, a learning rate of $52e-$ 5 and a warm-up ratio of $1\%$ are implemented. To enhance computational efficiency, the training process is conducted using distributed optimization with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) and is combined with Flash-Attention (Dao, 2024). Stage ➀ comprises $3$ epochs, whereas the satge ➁ and ➂ are conducted over $4$ epochs. A sequence length of $2,048$ tokens is used in Stages ➀ and ➁, which is increased to $4,096$ tokens in Stage ➂ to support more complex reasoning tasks. Furthermore, we employ an annealing strategy at the end of Stage ➂ with high-quality MPM samples, with the aim of further enhancing model accuracy in complex reasoning tasks.
A.4 Benchmarks
We report the statistics of the evaluation datasets in Table 6.
Table 6: The statistics of the evaluation datasets.
Dataset # of Test set Avg. Length
MATH 5,000 30.7
GSM8K 1,319 46.3
MiniF2f 244 30.5
A.5 Details of Metrics
Appendix B Additional Analysis
B.1 The Risk of Data Leakage
@@
B.2 Different Evaluation Strategies on Arithmetic Benchmarks
@@
Model MATH GSM8k
Pass@1 Maj@8 Pass@1 Maj@8
Llama-3.1-8B 58.18 59.46 84.00 87.26
DeepSeek-Math-Base-7B 66.74 71.70 88.70 91.43

### v4

@@ 6 Ablation Study
Figure 4: An evaluation of the effectiveness of the PPT strategy. We present the zero-shot Pass@1 results on the MATH and GSM8k benchmarks across three cumulative stages of the PPT strategy. The results highlight the PPT strategy’s cumulative effectiveness, showing increased performance with each progressive stage.
6.1 Impact of Stages in PPT Method
We evaluate the PPT method using two base models, Llama-3.1-8B and DeepseekMath-7B-Base, on the MATH and GSM8K benchmarks at different PPT stages. As shown in Fig.˜4, the vanilla models exhibit limited performance. After stage ①, performance improves significantly, with Llama-3.1-Base gaining $47.9\%$ on MATH and $61.0\%$ on GSM8K, emphasizing the need for mathematical understanding and model warming. Stage ②, utilizing NLR and AR data, yields limited gains compared to stage ① due to pre-training data already containing relevant content. However, this does not imply that stage ② is unimportant, as it plays a critical role in CoR by reactivating and refining computational skills essential for collaborating with other reasoning paradigms. Both stages ① and ② focus on reactivating models for mathematical tasks. In stage ③, training with three paradigms further enhances performance, suggesting that rare or unseen paradigms improve reasoning abilities. The consistent performance gains with added paradigms in  Fig.˜4 robustly demonstrate the superiority of our multi-paradigm PPT approach over single- or fewer-paradigm fine-tuning of the same base model, yielding a more powerful, comprehensive mathematical reasoner.
Benchmark
[…]
SR $\rightarrow$ Summary
[…]
AR $\rightarrow$ Summary
MATH 49.9 66.7
GSM8K 84.2 88.7
@@ 6.1 Impact of Stages in PPT Method
Model GSM8k@1 MATH@1 AMC2023 (Maj@64) AIME2024 (Maj@64) miniF2F@128 DSM + NLR 33.9 27.8 14/40 1/30 - DSM + AR 75.8 37.6 10/40 0/30 - DSM + SR - - - - 44.3 DSM + CoR (ours) 88.7 66.7 34/40 12/30 52.9
Table 5: Zero-shot results of multi- and single-paradigm fine-tuning on DeepSeekMath-7B-Base (DSM) for mathematical Reasoning benchmarks. The best results are in bold.
Model Size MATH GSM8k miniF2F
Qwen2.5-Math (Base) 1.5B 34.0 39.3 0.0
7B 51.8 90.0 0.0
Qwen2.5-Math (CoR) 1.5B 57.6 84.5 51.6
7B 64.7 90.0 52.5
Llama-3.1 (Base) 8B 4.2 6.2 25.8
70B 16.8 20.5 22.5
Llama-3.1 (CoR) 8B 58.2 84.0 53.3
70B 70.7 90.0 56.2
@@
Table 6: Zero-shot Pass@1 results on the mathematical benchmarks across different model scales. The best scores are in bold.
6.2 Order of the Reasoning Paradigms
As shown in Table˜4, we investigate the impact of varying the sequence of reasoning paradigms in CoR-Math-7B during zero-shot arithmetic tasks, modifying only the prompt structure. Given that NLR closely aligns with the pre-training of LLMs, it serves as the first paradigm. The results indicate that SR before AR resulted in the highest accuracy, with $66.7\%$ on MATH, compared to $49.9\%$ for AR before SR. A possible explanation is that SR plays a crucial role in decomposing the problem into manageable sub-steps, thereby providing a structured foundation for subsequent AR. This demonstrates that paradigms in CoR are interdependent, not isolated, and performance hinges on how they are linked during reasoning, beyond merely mastering individual paradigms. Furthermore, positioning AR preceding the Summary boxes may facilitate a more coherent integration of the computational outcome into the final answer.
6.3 Impact of Model Scales
Table˜6 evaluates base model performance under the CoR framework with varying parameters, such as Qwen2.5-Math models (1.51.5B / 77B) and Llama-3.1-8B (88B / 7070B). The results show that CoR scales with model size. The 77B Qwen2.5-Math outperforms the 1.51.5B version by $7.1\%$ on MATH and $5.5\%$ on GSM8K. Similarly, Llama-3.1 7070B gains $2.9\%$ over its 88B version on minF2F. This supports that CoR exhibits parameter scalability.
6.4 Multi- vs. Single-Paradigm Reasoning
To further validate the superior performance of multi-paradigm reasoning, we conducted an ablation experiment comparing it to single-paradigm fine-tuning We fine-tuned the DeepSeekMath-7B-Base model separately using single-paradigm paths (NLR, AR, SR) extracted from the MPM dataset. All experiments employed identical base models, problem sets, and training data sizes ($10,000$ samples per paradigm). Experimental results, summarized in Table 5, clearly demonstrate that our CoR model significantly outperforms single-paradigm fine-tuned models across diverse tasks. Specifically, for GSM8k and MATH datasets, CoR achieves improvements of $12.9\%$ and $29.1\%$, respectively, compared to the best-performing single-paradigm approaches. On AMC2023 and AIME2024, CoR solves $20$ and $11$ more problems, respectively, compared to single-paradigm approaches. Lastly, on minF2F, CoR achieves an accuracy improvement of $8.6\%$. Collectively, these improvements underline CoR’s effectiveness in integrating multiple reasoning paradigms, enabling robust and generalized problem-solving capabilities.
7 Conclusions
This paper introduces CoR, a novel unified reasoning framework that enhances mathematical reasoning in LLMs by synergistically integrating natural language, algorithmic, and symbolic. CoR tackles the limitations of single-paradigm approaches with the PPT strategy and the MPM dataset, enabling LLMs to progressively master diverse reasoning paradigms for improved generalization and performance. Our CoR-Math-7B outperformed SOTA models on five challenging benchmarks, showcasing enhanced zero-shot generalization. Ablation studies validated the benefits of progressive training and paradigm sequence. Additionally, CoR offers a new perspective on test-time scaling through multi-paradigm reasoning to enhance both efficiency and performance.
Acknowledgments
This work was partly supported by the Shenzhen Science and Technology Program ( JSGG20220831110203007 ) and the research grant No. CT20240905126002 of the Doubao Large Model Fund.
Limitations
This paper proposes a general zero-shot reasoning method that supports different paradigms to solve multiple tasks. Evaluating zero-shot performance is challenging, as many methods rely on a single evaluation metric, like zero-shot pass@1. Additionally, most approaches focus on few-shot settings, as seen in benchmarks like miniF2F, limiting the comprehensiveness of our evaluation with aligned settings. To address this, we include additional experiments in Sec.˜C.2 to examine the impact of different evaluation strategies. On the other term, CoR can explore a multi-paradigm solution space, covering up to three paradigms, such as in large-scale arithmetic tasks. However, due to resource constraints and the fact that other methods can only predict within a single paradigm, we did not further consider using SMPS on benchmarks such as MATH to ensure fairness.
Appendix
Appendix A Discussion on Reasoning Hierarchy: Paradigms, Paths, and Steps
@@
As shown in Fig.˜5, this paper posits that reasoning texts generated by LLMs exhibit a reasoning hierarchical structure, which consists of three levels: reasoning paradigms, reasoning paths, and reasoning steps.
@@
  • Reasoning paradigms comprise one or more reasoning paths. They often contain multiple potential final answers, thus necessitating a summarization method, such as a summary module, to extract the ultimate answer. Furthermore, a reasoning paradigm uses a single knowledge media, such as natural language.
Figure 5: An overview of the reasoning hierarchical structure. (a) A single reasoning paradigm depicts multiple distinct reasoning paths. (b) An example of single-paradigm reasoning includes one reasoning path, which contains some reasoning steps. (c) Multi-paradigm reasoning includes several distinct reasoning paradigms.
Figure 6: A comparison of reasoning methods in several advanced studies.
@@
Furthermore, current studies can be categorized based on their focus. As shown in Fig.˜5 (a), contemporary work concentrates within a single paradigm, optimizing along two dimensions: depth (the number of reasoning steps) and width (the number of reasoning paths). For instance, regarding reasoning depth, the CoT method (Wei et al., 2022) employs prompts to increase intermediate steps within one reasoning path to achieve higher performance. Concerning reasoning width, some approaches (Wang et al., 2023) involve altering sampling techniques to generate multiple distinct reasoning paths. Random sampling exemplifies this. Other studies (Zhu et al., 2023) employ scoring mechanisms for reasoning steps, guiding LLMs to generate several more desirable reasoning paths. Monte Carlo search illustrates this. Building upon these generated reasoning paths, existing work proposes various integration strategies to obtain the final answer, such as Best-of-N and Self-consistency (Wang et al., 2023). Specifically, as shown in Fig.˜6 (a), recent advanced studies introduce the deep reasoning method, which focuses on generating serial concatenation of reasoning paths followed by summarization (like OpenAI o1 OpenAI (2024)). Since these methods are based on different optimization dimensions, we can combine them feasibly in applications.
@@
Specifically, some approaches prioritize a dominant paradigm for reasoning yet hope to integrate other paradigms for guidance. For instance, as shown in Fig.˜6 (b), the reasoning process of ToRA (Gou et al., 2024) presents an interleaved reasoning method with different knowledge media. It involves initial natural language generation, which is followed by code generation. Subsequently, the process awaits the results of code execution before generating further natural language. InternLM2.5-StepProver (Wu et al., 2024) exhibits a similar pattern. It incorporates natural language annotations to support symbolic reasoning steps. However, these interleaved reasoning approaches do not constitute true multi-paradigm reasoning. The primary paradigm can independently achieve the final answer without the supplementary paradigms. Consequently, these methods imply a single-paradigm reasoning approach enhanced by other paradigms, rather than genuine multi-paradigm reasoning.
@@
Current methods often overlook the synergistic effects between paradigms and underestimate the importance of complete reasoning for achieving accurate results. To address this challenge, as shown in Fig.˜5 (c) and Fig.˜6 (c), we introduce Chain-of-Reasoning (CoR), a unified reasoning framework capable of multi-paradigm reasoning. This framework embodies the following potential advantages:
@@
Sequential reasoning dependency. As a type of multi-paradigm reasoning, CoR is not a mere accumulation of isolated steps; instead, it represents a coherent, interconnected process. The output from an earlier paradigm functions not only as input for a subsequent paradigm but also informs the foundational information for its reasoning. For instance, knowledge derived from a natural language paradigm guides subsequent algorithmic paradigms, thereby enhancing the efficiency and accuracy of code generation based on detailed natural language reasoning. This mechanism also provides substantial context from prior paradigms to later paradigms. It operates as a form of preliminary rehearsal where subsequent paradigms can follow correct reasoning steps or rectify incorrect ones.
@@
Appendix B Experiment Details
In this study, we use open-source tools, models, and datasets in compliance with their respective open-source licenses and solely for academic research purposes.
B.1 The detail of the Universal Text Template
@@
As shown in Fig.˜7, we apply the designed universal text template on all training samples.
B.2 Example Prompts for Dataset Enhancement
@@
Fig.˜8 and Fig.˜9 provide examples of such prompts designed for augmenting and refining formal proofs within the Lean 4 theorem prover for the Numina-TIR and Lean-Workbook datasets, respectively. The prompt specifies the input and output format and includes placeholders for the problem statement, informal proof, and corresponding formal proof in Lean 4. It also incorporates $N_{\text{shot}}$ few-shot examples to further guide the model. In our experiments, we follow the common settings (Jiang et al., 2023; Yang et al., 2024; Wang et al., 2025; Zhao and Zhang, 2024) and set $N_{\text{shot}}=8$. This structured approach and setting are applied to all datasets to ensure comprehensive coverage and logical consistency across the enhanced dataset.
B.3 Details of Training Settings
In all stages of the PPT method, a learning rate of 2e52e-5 and a warm-up ratio of $1\%$ are implemented. To enhance computational efficiency, the training process is conducted using distributed optimization with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) and is combined with Flash-Attention (Dao, 2024). Stage ➀ comprises $3$ epochs, whereas the stage ➁ and ➂ are conducted over $4$ epochs. A sequence length of $2,048$ tokens is used in Stages ➀ and ➁, which is increased to $4,096$ tokens in stage ➂ to support more complex reasoning tasks. Furthermore, we employ an annealing strategy at the end of stage ➂ with high-quality MPM samples, with the aim of further enhancing model accuracy in complex reasoning tasks.
B.4 Benchmarks
We report the statistics of the evaluation datasets in Table˜7.
Dataset # of Test set Avg. Length
MATH 5,000 30.7
GSM8K 1,319 46.3
AIME2024 30 59.2
AMC2023 40 47.2
MiniF2f 244 30.5
Table 7: The statistics of the evaluation datasets.
B.5 Details of Metrics
[…]
@@ B.5 Details of Metrics
B.6 Experimental Details of MATH and GSM8K benchmarks
For arithmetic-related results, we analyze the number of mathematical samples utilized by the baseline model following the pre-training phase, as illustrated in Fig.˜1. Specifically, as shown in Table˜8, we present the sample quantity along with the model performance.
Appendix C Additional Analysis
C.1 The Risk of Data Leakage
@@
Model MATH GSM8k Pass@1 Maj@8 Pass@1 Maj@8 Llama-3.1-8B 58.18 59.46 84.00 87.26 DeepSeekMath-Base-7B 66.74 71.70 88.70 91.43
@@
C.2 Different Evaluation Strategies on Arithmetic Benchmarks